Reports of 52766
Detector Operation (Operations Report)
gherardini - 22:55 Friday 25 November 2022 (57914) Print this report
Operator Report - Afternoon shift
No activity in this afternoon, just the recovery of SPRB multisas control done by Henk-Jan and some test on B1p_PD2 demodulation done by Alain; I left the north cavity locked.
Images attached to this report
AdV-DET (Commissioning)
gouaty, hui, masserot, mwas, nocera, casanueva, magazzu, tournefier - 19:06 Friday 25 November 2022 (57905) Print this report
Attempt to measure the SB amplitude noise in transmission of the OMC with RAMS on

The goal of this morning shift was to try to measure the side band amplitude noise, by locking the OMC on each side band with NI single bounce, and take a measurement of the B1 spectrum. This measurement was performed first on the 56 MHz without the RAMS, then it was performed on the 56 MHz, the 6 MHz and the 82 MHz with the RAMS activated.

Started OMC scan at 08h05 utc:

  • TEM00 carrier : 100 uW (power read on B1_PD3)
  • 6 MHz side band = peak at 2 uW, but we have still ~1 uW on carrier tail.
  • 56 MHz side band with about 0.4 uW

OMC locked on 56 MHz side band at 08h15m00 utc. Taking data until 09h00 utc.The obtained B1 spectrum is compared to the sensing noise on Fig.1.

At 10h00 utc, RAM servo is ON.

OMC locked on 56 MHz at 10h10m20 utc with 0.5 mW on B1, north arm locked. NE misaligned (single bounce) at 10h11m20 utc. B1_PD1/PD2 opened at 10h13m00 utc. Taking data for 5 min. Measurement shown on Fig.2 (reference plot in purple is the measurement taken previously without the RAMS). There is little difference between the two spectra.

56MHz SB reduced to 7 dB, but OMC unlocked.

8 MHz reduced by -16 dB at 10h41m16 utc.

OMC locked on 6 MHz at 10h48m00 utc with 2.4 mW on B1, but north arm is locked. B1_PD1/2 open at 10h49m40 utc Taking data for 3 min (Fig.3). The noise level is high, we suspect that this is because the north arm is locked.

North arm unlocked, NE misaligned to setup the single bounce, taking data from 10h53m30 utc (4 min). As shown on Fig.4, the noise level is better, but still quite high. We suspect that we are limited by noise from the carrier (half ot the power is from the carrier when OMC is locked on the 6 MHz).

Romain B is moving the SPRB lens at 10h57m45 utc. Taking data for 5 min (B1 spectrum seem to be slightly better). See Fig.5. OMC manually unlocked.

We put back the 8 MHz at full modulation depth at 11h05m32 utc.

OMC locked with B1_PD3 on 56 MHz at 11h15m40 utc. Hand off to B1_PD2 at 11h20m00 utc.

ITF in single bounce. Reduced 56 MHz SB by -7 dB at 11h25m19 utc.

Starting data taking with OMC locked on 56 MHz (with 0.1 mW on B1) at 11h25m30 utc. Taking data for 3 min30.

Putting back 56 MHz at 0dB at 11h29m04 utc. Taking data with 0.5 mW on B1 for 5 min. See Fig.6. There is no major difference compared to previous measurements on the 56 MHz.

Realigned OMC, OMC locked on 82 MHz in single bounce, with B1_PD2, at 13h37m40 utc (0.25 mW on B1). Taking data for 8 Min (Fig.7). We are not sure of the quality of the measurement because the 82 MHz is close to the carrier first order mode that we did not manage to reduce below 2%.

Trying to improve alignment while OMC locked on 82 MHz, but we are changing the shape of the beam. Aligning using a mode so close to the first order mode does not seem possible.

Scanning OMC from 14h00 to 14h22 utc (Fig.8): negative scan on Fig. 9 (with B1_PD1/2 opened, but triggering fast shutter at TEM00 crossing). Then during the positive scan B1_PD1/2 shutters were closed. To be analyzed later.

Fig.10 shows a smaller scan over the two first order modes. We note that we have an order 2 at 0.9 mW on B1 and an order 1 at 2.5 mW, which corresponds to a mode mismatch smaller than 1% and a misalignment defect smaller than 2%.

OMC locked again on 56 MHz at 14h56m30 utc (0.54 mW on B1). Taking data for 10 min (Fig.11). The noise level is unchanged.

We compare the B1 spectrum obtained with the OMC locked on the 56 MHz with the RAMS channels in Fig.12, 13, 14:

  • in purple when the RAMS was engaged at 15h00 utc.
  • in blue when the RAMS was off at 08h30 utc.

On these figures we observe that:

  • a clear reduction of noise on the RAMS channels when the RAMS is engaged.
  • a clear reduction of the coherence between the RAMS channels and the B1 spectrum, however some coherence remains.
  • there is actually a little improvement of the B1 spectrum between 40 and 90 Hz
Images attached to this report
AdV-PAY (NI and WI Payloads)
F.Travasso, H.Vocca - 18:42 Friday 25 November 2022 (57913) Print this report
WI monolithic suspension recovery
The WI mirror is in the cleanroom class 1 inside the box (Picture 1). The two magnets that felt during the crash leaving a small damage on the mirror (Picture 2) have been rebonded (Picture 3) and the residuals of the anchors at the level of the ears have been removed increasing the presence of holes (Pictures 4 and 5).
Images attached to this report
AdV-DET (Commissioning)
masserot - 18:15 Friday 25 November 2022 (57912) Print this report
B1p_PD2 demodulated at F25MHz (F81MHz-F56MHz)

For test purposes of the 25MHz (F81MHz-F56MHz) demodulation chain, the B1p_PD2 50MHz demodulation has been replaced by the 25MHz one .

As consequence the B1p_PD{1,2} demodulation phases may have to be tuned

Infrastructures (Air Conditioning)
R. Passaquieti - 17:52 Friday 25 November 2022 (57911) Print this report
Comment to New motor anda fan for NEB AHU (57814)
This afternon at about 14:10 UTC the 200 mm diameter pulley of the air fan inside the NEB AHU was substituted with a larger one of 250 mm diameter.
The inverter frequency driving the fan rate was increased accordingly from 28 Hz to 30 Hz.
Detector Operation (Operations Report)
amagazzu - 17:43 Friday 25 November 2022 (57908) Print this report
Operator Report - Morning shift

The shift was dedicated to the two DET Planned activities, B4 shape analysis with SPRB in vacuum (Bonnand, A. Daumas) and RAMS test by SB RIN measurement on B1 (Gouaty).
The first part of the shift was dedicated to venting the SPRB tower, carried out by the VAC Team from 7:40 UTC to 8:20 UTC. Afterward, we contacted the SBE expert (Bulten) that, from remote, closed the SPRB Loops: recovery concluded at 9:08 UTC. In the meantime, under the request of Gouaty, I set the ITF in NI Single Bounce (NARM_LOCK in MISALIGNED_NORTH_END) from 7:53 UTC.
From 9:00 UTC to 9:55 UTC Nocera, Casanueva and Sposito worked in the Electronic Injection Room to complete the implementation of the RAMS Servo, during this period the RFC was unstable and kept unlocking.
After this work was completed, I kept relocking the North Cavity and putting the ITF back in Single Bounce configuration under the request of Bonnand and Gouaty.
Activity on SPRB in Air completed at 13:50 UTC, the VAC team started the Evacuation of the minitower at 14:13 UTC. SPRB loops recovery started by Bulten at 15:00 UTC.
DET activity completed at 15:15 UTC, PR parked position restored and North Arm cavity left locked.

Other activities carried out during the shift:
- Installation of new Pulley for NEB UHA (Passaquieti, Tringali);

Sub-system reports

SBE
SPRB Position and Angular Loops opened at 13:50 UTC to allow the evacuation of the minitower.

TCS
Under request of Ballardin, at around 12:40 UTC I went to the NE Building to switch on the Motor Drivers related to TCS.

Images attached to this report
AdV-SGD (FDS commissioning)
vardaro - 17:16 Friday 25 November 2022 (57910) Print this report
New FDS long run started

I tuned again the parameter of FDS measurement:

  • increased the gain of GR longitudinal control Z_GAIN = 18 CC_GAIN = 4
  • improved the alignment in the FI isolator
  • tuned the new waveplate on EQB1
  • with the help on Henning retuned the OPA temperature

At 17:05 UTC we restarted the SQZ ON OFF measurement with 140 Hz of detuning

AdV-DET (Commissioning)
bonnand, daumas, magazzu - 16:45 Friday 25 November 2022 (57909) Print this report
Translation of SPRB_MMT_L2 lens with SPRB in air

This morning SPRB was vented by VAC team and Henk-Jan recontrolled the suspension and bench at around 9:30 LT.

At first we put the lens to the end of range on the "positive" side (meaning we asked for positive steps in the vpm).

Then we moved towards the end end of range by steps of -30 000 steps and at each step (up to step 11) we took data with the arm locked and unlocked.

After reaching the end of range in the "negative" direction, we moved the lens to the "positive" end of range again by steps of + 30 000 steps waiting about 1 minutes at each steps.

It is worth noting that the lens moves less in positive than in negative: about 13.5 steps of -30k steps to go from one end of range to the other end of range, whereas it is need about 20 steps +30k steps in the other direction.

We could observe that the beam is astigmatic also in air but with a different focusing than in vacuum.

Figure 1 shows the size of the beam on Cam1 and Cam2 during the whole activity.

Figure 2 shows the size of the beam on Cam1 and Cam2 during the way from "negative" to "positive" end of range.

The movement seems continuous but it is not, it just appears like this because the value given by the fit is fluctuating a lot in air.

Figure 3 shows the same as figure 2 but with the fit in X and Y superposed for the 2 cameras, one can see that at some point the value in X and Y are crossing for the two directions.

 

Some more analysis needs to be done but it will be more difficult quantitatively than in vacuum because of the fluctuation of the beam in air.

 

At the end of the range we increase the integration time and gain of Cam1 between 13:25 and 13:35 UTC to see more the foot of the beam where we see some pollution below the beam in GxC and on top in matlab images (figure 4).

Figure 5 and figure 6 show a vertical and horizontal slice of the beam.

This spot at the bottom is probably also there when the beam is larger vertically and might makes the beam larger in vertical and it could lead in some misinterpretation.

Images attached to this report
Non-image files attached to this report
AdV-EGO (Contamination Control)
menzione, ciardelli - 10:41 Friday 25 November 2022 (57906) Print this report
Contamination control inside CEB Clean Rooms

During the last few weeks we have been monitoring the contamination inside the Class100 Clean room of the Central Building.
Yesterday morning 11/24 at 9:00 UTC we moved the particle counter from the Class100 to the Class1 Clean room to verify that the environment was indeed in the perfect conditions to accommodate the WI mirror. As expected the contamination inside the Class 1 flux hood is 0. Trend attached

Images attached to this report
AdV-PSL (Commissioning of SS)
mwas - 9:34 Friday 25 November 2022 (57904) Print this report
PSTAB 50Hz harmonics measurement

The OMC lock on single bounce beam provides some interesting check of how the laser power stabilization is performing.

Figure 1 shows the spectrum of the B1 photodiodes, PSTAB PDa (in-loop?), PSTAB PDd (out-of-loop?), and RFC TRA. On B1 there are many bumps likely due scattered light from the spurious beams from misaligned PR and SR. But the noise floor at ~1kHz is relatively coherent with PSTAB signals. There is also a large amount of 50Hz harmonics visible on B1 up to a few kHz, and their relative height looks very similar to what is visible on B1

Figure 2 shows the B1 spectrum normalized to get the RIN. The noise floor is rather high at 7e-9 1/rtHz, while from what I remember the PSTAB specification is 2.5e-9 1/rtHz. How does one calibrated the PSTAB photodiodes into RIN?

The level of 50Hz harmonics appear a factor few higher than what it was at the end of O3.

 

 

Images attached to this report
Detector Operation (Operations Report)
amagazzu - 8:22 Friday 25 November 2022 (57903) Print this report
Operator Report - Initial Status

ITF Mode: NOT_LOCKED;
ITF State: North cavity locked;
Quick Summary: All Suspension Loops closed with the exception of WI controls; SPRB Position and Angular loops opened by Bonnand at 6:56 UTC, FCEM actuator2 below the threshold, all other SBEs loops closed; WI tower open;
Activities ongoing since this morning: No known activity ongoing.

Images attached to this report
Detector Operation (Operations Report)
berni - 22:55 Thursday 24 November 2022 (57901) Print this report
Operator Report - Afternoon shift

The activities of the shift were the following:

At 15:25 UTC the north cavity was locked and then at 15:37 UTC the ITF was put in single bounce. At 16:15 UTC the north cavity was locked; still locked at the end of the shift.

 

Note about PR position:

+40urad TX and  -40urad TY from the standard parking position.

Images attached to this report
AdV-PAY (NI and WI Payloads)
travasso, majorana - 21:50 Thursday 24 November 2022 (57900) Print this report
Comment to Extraction of WI payload and recovery preparation (57896)
Status of the WI payload
Images attached to this comment
Infrastructures (Air Conditioning)
Tringali, Fiori, Paoletti, Passaquieti - 18:44 Thursday 24 November 2022 (57866) Print this report
Comment to New motor anda fan for NEB AHU (57814)

The installation of the new fan with a 200 mm pulley produced a significant acoustic broad-band noise reduction in the experimental hall, Figure 1 and Figure 2 (red curve). A seismic noise reduction is observed in narrow frequency regions, Figure 3 and Figure 4 (red curve).
The same noise reduction (blue curves) is visible in the accelerometers installed on the AHU box (*UTA*BOX), supply and return air ducts, and in the microphones of the experimental all (NN*INF_02/03/04), SAS room (*SAS_MIC and NN*INF_05) and AHU room (NN*INF_06), Figure 5 and Figure 6.
The environmental parameters of the experimental hall (pressure, temperature, humidity) are stable after the last slow-down of the inverter (28 Hz), Figure 7.
With the inverter set to 28 Hz, the frequency values of the motor-fan-belt system are:

motor ~ 13.95 Hz
fan  ~ 10.47 Hz
belt ~ 4.74 Hz

Images attached to this comment
AdV-INJ (Input Mode Cleaner cavity)
derossi, gosselin, sposito, spinicelli - 18:14 Thursday 24 November 2022 (57899) Print this report
EOM_corr faulty cable replaced

This morning, during some investigations on the FModErr signal in the "piscina" we noticed that the IMC could lock anymore. Looking closely at the rampauto cabling, we found that the EOM_corr BNC cable was broken at the connection to the device (the connector was separated from the signal wire), thus preventing the rampauto to send the correction.
Once the connector has been replaced, the IMC relocked immediatly.

Images attached to this report
AdV-INJ (Beam Pointing Control)
derossi - 18:07 Thursday 24 November 2022 (57898) Print this report
BPC noise injections

We injected noise on the BPC piezo actuators, with the IMC automatic alignment disabled (only in drift control).

The gps start are: TX = 1353340973, TY = 1353341493, X = 1353342038, Y = 1353342514, the duration 300 s for each injections (plots 1 to 4).

Figures 5 to 8 show the transfer functions with the RF quadrants used for the IMC automatic alignment tfestimate(Bsx TX/TY/X/Y , INJ IMC QD NF/FF I H/V)

Images attached to this report
AdV-PSL (Commissioning of SS)
derossi, gosselin, montanari, spinicelli, sposito - 18:02 Thursday 24 November 2022 (57897) Print this report
Comment to AdV+ PStab --air side (57820)

Today we added a cable from SIB1 platform to the "piscina"  in order to be able to use the new PSTAB_PDb_DC as a trigger for the IMC rampauto as it was before.

 

AdV-INJ (Input Mode Cleaner cavity)
gosselin, derossi, spinicelli, sposito, montanari - 17:49 Thursday 24 November 2022 (57893) Print this report
Investigation on the DAQ_FMOD_Out0 RF pollution

This morning we investigated on the RF pollution seen by the signal INJ_IMC_TRA (cf entry https://logbook.virgo-gw.eu/virgo/?r=57073)
For the RF frequency band one should look at the signal DAQ_FMOD_Out0_sample, which is the one used for the digital Fmoderr.

In the previous entry we noticed that even when there is no beam arriving on the photodiode, there were anyway lines at 6 8 56 MHz.

We did some tests on the cables bringing the signal from the EIB towards the demodualtion box.
Here below is the tests that we did, times are in local time :

10:55:00: 50 Ohm at the level of the lemo/BNC connection on EIB 
10:58:00 same but without the 50 Ohm
11:08:50 replace the 15m-long LEMO cable from EIB by a BNC cable
11:14:00 BNC cable and 50 ohm
11:16:20 Ferrite block on the BNC cable at the EIB level
11:19:00 Ferrite block on BNC cable in piscina
11:20:15 remove the ferrite block at the EIB level
11:22:15 put a second ferrite block on the BNC cable in piscina
11:xx:xx switch back to the LEMO with now a magnet on the piscina side
11:35:00 With LEMO unplugged from the BNC/LEMO box in piscina (ie only box and few meters of BNC towards the demodulation box)
11:38:40 BNC unplugged from the box
11:41:00 BNC plugged to box without other splitter
11:44:40 BNC plugged without LEMO
11:48:20 we try to connect LEMO to BNC with another LEMO/BNC box
11:52:50 we now plug the 15m long BNC directly to BNC in piscina (with an I)
12:00:30 we change the position of the RF splitter in order to separate the interesting signal befrore arriving on the T splitter
12:02:00 we come back to LEMO-BNC box with the RF splitter

 

Some conclusion from thoses tests :

- the 8 MHz disappeared. To be noted however that amplitude is set to 0 dB while the others are set to 12 dB
- best configuration BNC to BNC for the 56 MHz (11:52:50 configuration), lemo to BNC for the 6 MHz (11:41:00) 
- Strangely using a dedicated lemo to BNC box worsen the situation
- the "signal distribution", ie 2 BNC splitters at the output of the lemo to BNC box, are worsening the situation
- the 2f lines that can be seen in DAQ_FMOD seem to come from the optics. The 112 MHz line was reduced by two order of magnitude while unplugging the first cable (10:55:00) and the 12 MHz one disappear. 

The origine of the pollution is now better understood. We should speak with some expert to try to get rid of it.

AdV-PAY (NI and WI Payloads)
Travasso, Menzione, Ciardelli, Pasqualetti , Majorana - 17:43 Thursday 24 November 2022 (57896) Print this report
Extraction of WI payload and recovery preparation
AT 11:45 we entered in the WI vacuum chamber and at 14:00 the payload was down on the HYMO. Then, the payload was transported in the clean room. Flavio is preparing the case for the mirror. We have the plan to settle the mirror in CL1 in the late evening or tomorrow morning.
M1 mirror of TCS HS as a whole and we are confident to recover an acceptable prealignment. We followed the foreseen procedure.
No particular problem has been encountered.
Comments to this report:
travasso, majorana - 21:50 Thursday 24 November 2022 (57900) Print this report
Status of the WI payload
Images attached to this comment
AdV-DAQ (Calibration)
lagabbe - 17:35 Thursday 24 November 2022 (57895) Print this report
Comment to PCal photodiode timing measurement (57887)

The model of the ADC channel consists in the product of:

  • a low pass butterworth filter, 5th order, with the cut off frequency 80 kHz,
  • a low pass butterworth filter, 8th order, with the cut off frequency 7499 Hz,
  • an advance of 49 µs,
  • a residual delay, which is a parameter to be fitted to the measured frequency response.

The frequency response of Tx_PD1, is measured by computing the transfers function from the excitation signal to the Tx_PD1 signal, when the LED is flashing.
The frequency response of Tx_PD2 and Rx_sphere, is measured by computing the transfers function from the Tx_PD1 signal to the Tx_PD2 or Rx_sphere signal, when the white noise is inected in laser command signal.
Values of the residual delays between the model and the measure of the frequency response of the photodiodes.

ref signal test signal delay between ref and test signals (ns) residual delay of test signal (ns)
excitation NE_Tx_PD1 699 +/- 32 699 +/- 32
NE_Tx_PD1 NE_Tx_PD2 -79 +/- 1 620 +/- 32
NE_Tx_PD1 NE_Rx_sphere 91 +/- 6 790 +/- 32
excitation WE_Tx_PD1 649 +/- 36 649 +/- 36
WE_Tx_PD1 WE_Tx_PD2 -44 +/- 1 605 +/- 36
WE_Tx_PD1 WE_Rx_sphere 879 +/- 6 1528 +/- 36

The WE_Rx sphere has a follower circuit at its output, and WE_Rx has not. The follower circuit add a delay in the frequency response.

 

Images attached to this comment
AdV-TCS (Wavefront sensing)
nardecchia - 16:50 Thursday 24 November 2022 (57894) Print this report
HWS-DET beam check

As requested by Fiodor, the HWS-DET beam reflected back by the NI has been checked.

After switching OFF the HWS-INJ SLED, turning ON the HWS-DET SLED and opening the flip mirror on EDB, the beam appeared on the  HWS-DET CCD (at around 15.40 UTC).

 

Images attached to this report
Detector Operation (Operations Report)
gherardini - 15:59 Thursday 24 November 2022 (57884) Print this report
Operator Report - Morning shift

The main activities of this morning:

- fmodErr noise investigation;
- west input payload dismounting from the tower;

note: PR position: +40urad TX -40urad TY from the standard parking position.

Sub-system reports

SBE
recovery of SPRB vertical correction with F0 step motors from 8:30UTC to 8:40UTC.

Images attached to this report
Infrastructures (Air Conditioning)
R. Passaquieti - 15:10 Thursday 24 November 2022 (57892) Print this report
Comment to New motor anda fan for NEB AHU (57814)
I forgot to mention that yesterday the complete NEB AHU Inverter frequency tuning was from 32 Hz --> 30 Hz --> 28 Hz
Optical characterization (Optical characterization)
gherardini, sorrentino - 15:08 Thursday 24 November 2022 (57891) Print this report
B1p phase camera in single bounce vs PR mirror parking position

This morning we briefly investigated the impact of PR angular position on the fringe pattern that is visible on the B1p phase camera in single bounce configuration.

Fisrt picture: PR TX position vs time

Second plot: B1p phase camera images in original PR position

Third plot: B1p phase camera images after 20 urad tilt: fringes are rotated

Fourth plot: B1p phase camera images after 100 urad more tilt: fringes are no more there

Images attached to this report
AdV-DAQ (Data Acquisition and Global Control)
masserot - 14:38 Thursday 24 November 2022 (57890) Print this report
SSFS dbox resynchronisation

I found the SSFS_timing_error at 10ns instead of 0 : the jump occured at  20221124-9h46-UTC. So the SSFS DBOX has been reconfigured  to recover the correct running conditions

Operation perfomed at  2022-11-24-13h24m51-UTC    Reconfigure CEB_DBOX_SSFS in SSFS_dbox

 

Search Help
×

Warning

×